OpenStudy (anonymous):

Which mixed number falls between three and two-thirds and four and five-sixths?

hero (hero):

I'm almost certain there are a list of answer choices to choose from.

OpenStudy (anonymous):

YES THERE ARE

hero (hero):

There are many mixed numbers that can fit between those two

hero (hero):

So you should provide the list of answer choices

OpenStudy (anonymous):

Which mixed number falls between three and two-thirds and four and five-sixths? four and eleven-twelfths three and three-eighths four and one-fourth three and one-half

OpenStudy (anonymous):

U GOT IT

hero (hero):

It's pretty much the same thing we did last time, except a little more common sense applies.

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

SO WHAT DO WE DO

hero (hero):

I'll try my best to explain. But maybe I should explain in vyew.

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4e2458b70b8b3d38d3b71ab5 if u are seeing here u can understand the method

hero (hero):

There's a much easier method than that.

hero (hero):

It's just difficult to show on here

OpenStudy (anonymous):

OKAY

OpenStudy (anonymous):

3 2/3 = 12/3 4 5/6 = 29/6 4 11/12 = 59/12 3 3/8 = 27/8 4 1/4 = 17/4 3 1/2 = 7/2

hero (hero):

That's not the easy way either. If anyone wants to know the easier way, come here: http://vyew.com/room#/418625/Untitled_Room

OpenStudy (anonymous):

find least common denominator of the 2 given numbers and one of the multiple choice numbers. then you should be able to tell which one fits

hero (hero):

Nope, that's not the easier way either.

hero (hero):

It's the cross mutliplication technique but you can only do it with mixed numbers that have the same whole number

hero (hero):

So you can only compare 3 and 2/3 with 3 and 3/8 or with 3 and 1/2 for example

hero (hero):

And you can only compare 4 and 5/6 with 4 and 11/12 or 4 and 1/4

hero (hero):

You just have to make sure that whatever the mixed number is, it is between 3 and 2/3 and 4 and 5/6

hero (hero):

@Vivi305, you already know the cross multiplication technique.

OpenStudy (anonymous):

SO WAT IS THE ANSWER

hero (hero):

All you have to do is compare the fractions of the ones with the same whole number

OpenStudy (anonymous):

NVM

hero (hero):

Pretty simple

hero (hero):

Okay, I'll show you, but it is going to take a lot of work for me since you can't hear me on vyew.

hero (hero):

So we want to see which number falls between 3 and 2/3 <---> 4 and 5/6. So first, let's try 4 and 11/12  Since 66 > 60, that can't be it.